# Correcting Annotator Bias in Training Data: Population-Aligned Instance Replication (PAIR)

### v1

@@ Abstract
Models trained on crowdsourced labels may not reflect broader population views when annotator pools are not representative. Since collecting representative labels is challenging, we propose Population-Aligned Instance Replication (PAIR), a method to address this bias through statistical adjustment. Using a simulation study of hate speech and offensive language detection, we create two types of annotators with different labeling tendencies and generate datasets with varying proportions of the types. Models trained on unbalanced annotator pools show poor calibration compared to those trained on representative data. However, PAIR, which duplicates labels from underrepresented annotator groups to match population proportions, significantly reduces bias without requiring new data collection. These results suggest statistical techniques from survey research can help align model training with target populations even when representative annotator pools are unavailable. We conclude with three practical recommendations for improving training data quality.  [^1] The code for experiments will be available at https://github.com/soda-lmu/PAIR.
@@ 1 Introduction and Inspiration
NLP models should align with the interests and judgments of the population they impact Sorensen et al. (2024); Fleisig et al. (2024). However, the training and feedback data for these models often comes from crowdworkers or convenience samples of annotators (e.g.  student assistants). These populations differ from the general population on important characteristics like age, education, and cultural context Smart et al. (2024); Berinsky et al. (2012); Ouyang et al. (2022), and these characteristics impact the labels they assign Sap et al. (2022); Fleisig et al. (2023); Kirk et al. (2024).
@@
Fortunately, survey researchers have developed robust statistical techniques to estimate population-level parameters from non-representative samples Eckman et al. (2024); Bethlehem et al. (2011a). The top panel of Figure 1 shows a simple workflow of collecting survey data and then creating statistical weights to match the data to the population. We propose that similar techniques could help align machine learning models with target populations, even when working with imperfect annotator pools (bottom panel).
To test this approach, we use a simulation study, a common approach in the statistics literature (Burton et al., 2006; Valliant, 2019; Morris et al., 2019). We simulate seven populations, create several labeled datasets with varying mixes of annotators, and train model on each dataset. We investigate two research questions: RQ1: How does the composition of the annotator pool impact model calibration and performance? RQ2: Can techniques from survey methodology mitigate the effects of non-representative annotator pools?
@@
Our results demonstrate that models trained on nonrepresentative annotator pools perform poorly. However, simple adjustment methods can improve performance without collecting additional data. These findings suggest that insights from survey methodology help make AI systems more representative of the populations they serve.
2 Annotation Simulation
To address our research questions, we create populations with two types of people: those more likely to perceive offensive language and hate speech and those less likely. We then simulate labels, varying the mix of the two types of people, to approximate when happens when the pool of annotators does not represent the characteristics of the population.
Data.
We use a dataset [^2] https://huggingface.co/datasets/soda-lmu/tweet-annotation-sensitivity-2 of 3,000 tweets sampled from Davidson et al. (2017). Each tweet has 15 annotations of both offensive language (OL: yes/no) and hate speech (HS: yes/no) (Kern et al., 2023). We chose this dataset because the high number of annotations of each tweet gives us a diverse set of labels to work with.
We randomly select (without replacement) 12 labels (of both OL and HS) from the 15 labels of each tweet in the original dataset. [^3] For reasons that will become clear, it is helpful for the number of labels per tweet to be divisible by four. Let $p_{i,OL}$ be the proportion of the 12 annotators who labeled tweet $i$ as OL and $p_{i,HS}$ defined similarly. Table 1 shows the distribution of these proportions.
Variable p25 Median Mean p75
@@ 2 Annotation Simulation
p refers to percentile
Table 1: Distribution of $p_{i,OL}$ and $p_{i,HS}$ in Gold Dataset
Population Setup.
We imagine populations made up of equal shares of two types of people. Type A people are less likely to say a tweet contains OL or HS. Type B people are more likely:
@@
The HS probabilities are defined similarly. Here $\beta$ captures the magnitude of the bias. We vary $\beta$ from $[0.05,0.3]$ by $0.05$ to create seven populations and seven vectors of probabilities ($p_{i,OL}^{A},p_{i,OL}^{B},p_{i,HS}^{A},p_{i,HS}^{B}$) for each tweet.
Label Simulation.
For each value of $\beta$, we create four datasets, each with 3,000 tweets (Table 2). The Balanced Dataset contains OL labels from six A annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{A})$) and six B annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{B})$). The proportion of A and B annotators in this dataset matches the population. The labels in this dataset are our gold standard.
@@
Balanced 12 6 6
Unbalanced 1 9 6 3
Unbalanced 2 12 9 3
@@
NA = Not Applicable; * 3 B labels duplicated
Table 2: Four Training Datasets for each label (OL, HS) and each bias value ($\beta$)
We then create two unbalanced datasets. Unbalanced 1 randomly deletes three B labels for each tweet from the Balanced Dataset. Unbalanced 2 adds three additional A labels, drawn from $p_{i}^{A}$, to the Unbalanced 1 dataset. The Unbalanced 2 Dataset is more unbalanced than Unbalanced 1, but contains the same number of annotations as the Balanced dataset.
Finally, we create the Adjusted Dataset. It is the same as the Unbalanced 1 dataset, but the B annotations are duplicated. This duplication is an easy way to adjust the unbalanced training dataset to reflect the population; a generalized version of this adjustment is provided in A.5.
Appendix Figure 5 shows the percentage of tweets labeled OL and HS in the four datasets for each value of $\beta$.
3 Model Training and Evaluation
@@
We train models on each dataset. We divide each dataset, at the tweet level, into training (2000 tweets), development (500), and test (500) sets.
@@
We used RoBERTa base Liu et al. (2019) as our text classifier. The model trained for five epochs on each dataset, with development set optimization. To ensure reliable results, we trained five versions with different random seeds and averaged their performance. Appendix §A.2 contains additional details.
@@
We measure model performance using two metrics.
  • Absolute Calibration Bias (ACB): This metric compares the observed proportion of OL/HS labels in the Balanced dataset to each the model’s predicted probabilities ($\text{preds}_{i}$):
@@
    ACB simplifies the ECE metric Naeini et al. (2015); Guo et al. (2017) by removing bins, and makes the single difference more straightforward by incorporating true frequencies directly. A low ACB score means the model’s predicted probabilities match the true frequencies in the Gold dataset.
  • F1 Score: This standard classification metric balances precision and recall.
The appendix also presents ECE (similar to ACB) and ROC-AUC (similar to F1) metrics (see §A.4). We report all metrics on the test set.
@@ 4 Results
Model Calibrations.
Figure 2 compares the ACB in the test set for models trained on the simulated datasets. Lower ACB indicates better model calibrations with the test labels. The dark lines show average ACB across the five training runs and the shading shows the standard deviation. See Appendix Figure 7 for the ECE results, which are similar.
In the OL graph, ACB for the models trained on the Balanced and Adjusted datasets do not increase with $\beta$ and are close together. ACB for the models trained on the two unbalanced datasets is greater and grows with $\beta$. These results demonstrate the effectiveness of our adjustment method. Duplicating the labels from the underrepresented annotator type to match population proportions improves calibration.
In the HS graph, the trends are less clear. Because HS is rarer in the dataset (Table 1), the A annotations are often 0, which complicates interpretation of the HS panel of Figure 2. We address this issue with additional analyses below.
Figure 2: Model ACB scores, by dataset and bias ($\beta$)
Model Predictions.
Figure 3 compares the models’ F1 scores. In contrast to Figure 2, we do not see strong differences between the models trained on the different datasets. For all datasets, model performance declines with $\beta$: as the amount of noise in the labels increases, the models have a harder time predicting the labels. The ROC-AUC graphs show the same pattern (Appendix Figure 8).
Because the F1 metric focuses on binary predictions, it is less sensitive to training biases compared to calibration metrics, which more explicitly capture biases through prediction scores. These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models, and binary classification performance alone can obscure such effects. In decision-making, miscalibrated predictions can have harmful consequences when, e.g., hateful content remains undetected Van Calster et al. (2019).
Figure 3: Model F1 scores, by dataset and bias ($\beta$)
Results with Difficult Tweets.
Our simulations assumed that all tweets are impacted the same way (Eq. (2)), which is an oversimplification. More realistically, annotator characteristics likely have more impact for ambiguous tweets. We repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$. This approach not only focuses on those tweets where annotator characteristics likely play a larger role, it also eliminates the floor and ceiling effects in Eq. (2)). The filtered datasets contain 267 (OL) and 360 (HS) tweets.
Figure 4 shows the ACB for models trained on the filtered datasets. In the OL graph, the patterns in the ACB scores are similar to Figure 2. The Balanced and Adjusted models have similar ACB and are lower than the Unbalanced models. The HS graph shows the same pattern, which is in contrast to Figure 2 where HS performed differently.
The F1 results for this subset of tweets are in Appendix Figure 6. The OL graph shows higher F1 scores for the Balanced and Adjusted models than for the two unbalanced models. There is no clear pattern in the F1 scores for the F1 model, though we do see an unexpected decrease in agreement in the Adjusted model when $\beta=0.3$.
Figure 4: Model ACB scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$), by dataset and bias ($\beta$)
@@ 5 Discussion & Recommendations
Our results demonstrate two key findings about annotator representation in training data. First, models perform less well when trained on data from non-representative annotator pools. Second, simple statistical adjustments can help correct for these biases without collecting additional data. These findings have important implications for dataset creation and model training.
Recommendations
If these findings hold up in future work, we recommend the AI/ML field take three steps:
1) Identify which annotator characteristics influence labeling and feedback decisions for different tasks. Research should integrate social science methods to understand how demographics, attitudes, and behaviors shape annotation patterns Eckman et al. (2024). This knowledge would help determine which characteristics need to be balanced across annotator pools for specific tasks.
2) Begin collecting relevant characteristics from annotators and gather corresponding population-level data. National censuses and large-scale surveys could provide useful population benchmarks. [^4] Collection and release of annotator characteristics or weights derived from them may raise concerns about annotator confidentiality. This topic is outside the scope of this paper, however, the survey literature contains useful approaches (see Karr, 2016, for a review). We also note that collecting annotator characteristics may require involvement of Institutional Review Boards or other participant protection organizations Kaushik et al. (2024).
3) Implement data adjustment methods that account for differences between annotator and population characteristics. While our simple duplication approach showed promise, more sophisticated statistical techniques from survey research may yield better results.
@@
Stylized Biases.
@@
We have created only one version the four datasets for each label type and value of $\beta$, each of which contains random draws from the Bernoulli distribution. A more traditional statistical approach would create multiple versions of the datasets and train models on each one, to average over the sampling variability. We have not done that in this preliminary study because of the high cost and time needed to fine tune many RoBERTa models. As discussed, we have used five seeds in model training.
@@
While we measured calibration and classification accuracy, we did not examine other important metrics such as fairness across subgroups or robustness to adversarial examples. Future work should on training data adjustment should assess a broader range of performance measures.
@@ Acknowledgments
Appendix A Appendix
A.1 Data
Figure 5 shows the percentage of tweets labeled OL and HS in the four datasets for each value of $\beta$. The percentage in the Adjusted dataset is similar to that in the Balanced dataset for all values of $\beta$. The two unbalanced datasets have lower rates of OL and HS, because they overrepresent the A annotators, who are less likely to label OL and HS.
Figure 5: Percentage of OL, HS “yes” labels, by dataset and bias ($\beta$)
Because HS is rare in our dataset, as $\beta$ increases, a higher proportion of $p_{i,HS}^{A}$ are 00 while the $p_{i,HS}^{B}$ values increase. This issue leads to a higher proportion of “yes” HS labels in the Balanced and Adjusted datasets, which have more B labels than the unadjusted datasets. [^5] If we had picked the B labels to overrepresent, we would have seen a converse problem with the OL labels.
A.2 Model Training Details

### v2

@@ Abstract
Models trained on crowdsourced labels may not reflect broader population views, because those who work as annotators do not represent the population. We propose Population-Aligned Instance Replication (PAIR), a method to address bias caused by non-representative annotator pools. Using a simulation study of offensive language and hate speech, we create two types of annotators with different labeling tendencies and generate datasets with varying proportions of the types. We observe that models trained on unbalanced annotator pools show poor calibration compared to those trained on representative data. By duplicating labels from underrepresented annotator groups to match population proportions, PAIR reduces bias without collecting additional annotations. These results suggest that statistical techniques from survey research can improve model performance. We conclude with practical recommendations for improving the representativity of training data and model performance.  [^1] The code for experiments is available at https://github.com/soda-lmu/PAIR.
@@ 1 Introduction and Inspiration
When a hate speech detection model flags harmless expressions as toxic, or a content moderation system fails to identify genuinely harmful content, the root cause often lies not in the model architecture, but in who labeled the training data. While NLP models aim to serve broad populations, the human judgments used to train these systems typically come from a non-representative pool of annotators – crowdworkers and convenience samples whose demographics, cultural contexts, and worldviews may differ from the communities the models ultimately impact Sorensen et al. (2024); Fleisig et al. (2024). These non-representative annotator pools can have real consequences, because annotator characteristics like age, education level, and cultural background impact how content is labeled Sap et al. (2022); Fleisig et al. (2023); Kirk et al. (2024). Models trained on non-representative data can perpetuate the biases and blind spots of their training data Smart et al. (2024); Berinsky et al. (2012); Ouyang et al. (2022); Mehrabi et al. (2021); Rolf et al. (2021); Favier et al. (2023); Hebert-Johnson et al. (2018); Hüllermeier and Waegeman (2021).
@@
Fortunately, survey researchers have developed robust statistical techniques to produce population-level estimates from non-representative samples Bethlehem et al. (2011). The top panel of Figure 1 shows a simple survey workflow: collecting survey data, creating statistical weights to match the sample to the population, and estimating population parameters. We propose adapting this approach to the machine learning context, enabling models to better align with target populations even when trained on non-representative annotator pools (bottom panel).
To test this approach, we conduct a rigorous simulation study following established practices in statistical research (Burton et al., 2006; Valliant, 2019; Morris et al., 2019). Using controlled experiments, simulated populations, and multiple training datasets varying in annotator composition, we investigate two questions:
  • RQ1: How do non-representative annotator pools impact model calibration and accuracy?
  • RQ2: Can survey techniques effectively mitigate these annotator pool effects?
@@
Our results demonstrate that models trained on nonrepresentative annotator pools perform poorly. However, simple adjustment methods can improve performance without collecting additional data. These findings suggest that insights from survey methodology can help AI systems better represent the populations they serve.
2 Related Work
Several strands of related work inform our approach to identifying and mitigating bias due to the use of non-representative annotators:
Annotator Impact on Data and Models. Annotator characteristics and attitudes significantly influence label quality, particularly for subjective tasks like toxicity detection (Giorgi et al., 2024; Prabhakaran et al., 2021; Fleisig et al., 2023; Sap et al., 2022). For example, annotators’ political views and racial attitudes affect their toxicity judgments (Sap et al., 2022). Models trained on non-representative annotator pools inherit these biases and generalize poorly (Smart et al., 2024; Berinsky et al., 2012; Ouyang et al., 2022; Mehrabi et al., 2021; Rolf et al., 2021; Favier et al., 2023).
Annotator Demographics. Several researchers advocate collecting annotator demographics to assess representation and identify biases (Bender and Friedman, 2018; Prabhakaran et al., 2021). [^2] In our context, these characteristics are used only to analyze bias; because they are not available for unlabeled text, they are not features that the model can use. However, collecting and releasing these data can raise privacy concerns (Fleisig et al., 2023).
Debiasing & Data Augmentation Methods. Prior work has proposed various approaches to reduce bias in training data features and labels (Calders et al., 2009; Kamiran and Calders, 2012; Sharma et al., 2020). Most similar to our work is the resampling and reweighting approaches of Calders et al. (2009) and Kamiran and Calders (2012), and the oversampling of minority class cases of Ling and Li (1998). PAIR adapts these methods to balance annotator characteristics rather than class labels or sensitive observation-level features. PAIR retains the simplicity and interpretability of earlier resampling methods while extending them to a “Learning with Disagreement” (Uma et al., 2021; Leonardelli et al., 2023) setting with multiple annotations per observation, by replicating labels from underrepresented annotator groups.
3 Annotation Simulation and Model Training
To address our research questions, we imagine a population made up of two types of people: those more likely to perceive offensive language and hate speech and those less likely. We create three sets of simulated annotations which differ in the mix of the annotator types. We then create a fourth data set, using the PAIR algorithm, to attempt to fix the imbalance in the annotators. We fine-tune RoBERTa models on the four data sets and evaluate the effect of annotator characteristics on model performance (RQ1) and the ability of the PAIR algorithm to improve performance (RQ2).
3.1 Simulating Annotations
We use a dataset of 3,000 English-language tweets, each with 15 annotations of both offensive language (OL: yes/no) and hate speech (HS: yes/no) (Kern et al., 2023). [^3] https://huggingface.co/datasets/soda-lmu/tweet-annotation-sensitivity-2 We chose this dataset because the high number of annotations of each tweet gives us a diverse set of labels to work with. We randomly select (without replacement) 12 labels (of both OL and HS) of each tweet in the original dataset. [^4] As shown in Table 2, we can more carefully control the construction of our data sets when the number of labels per tweet is divisible by four. Let $p_{i,OL}$ be the proportion of the 12 annotators who labeled tweet $i$ as OL and $p_{i,HS}$ defined similarly. Table 1 shows the distribution of these proportions across the 3,000 tweets. The HS labels are unevenly distributed, whereas the OL labels are relatively balanced. Because this work is a preliminary investigation of the PAIR approach, and balanced labels provide a clearer view of PAIR’s abilities, the main body of the paper focuses on results with the OL labels; the HS results in Appendix §C.
Variable $25^{th}$ pctile. Median Mean $75^{th}$ pctile.
@@ 3.1 Simulating Annotations
Table 1: Distribution of $p_{i,OL}$ and $p_{i,HS}$ in original data
We imagine a population made up of equal shares of two types of people. Type A people are less likely to say a tweet contains OL. Type B people are more likely:
@@
Here $\beta$ captures the magnitude of the bias. We vary $\beta$ from $[0.05,0.3]$ by $0.05$, corresponding to an increase or decrease in the probability to judge a tweet as OL by five to 30 percentage points. This range is large on the probability scale and covers most reasonable situations. With these seven values of $\beta$, we create seven vectors of probabilities ($p_{i,OL}^{A},p_{i,OL}^{B}$) for each tweet.
We then create four datasets, each with 3,000 tweets (Table 2), for each value of $\beta$. The Representative Dataset contains OL labels from six A annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{A})$) and six B annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{B})$). Because the proportion of A and B annotators in this dataset matches the population, the labels in this dataset are our gold standard.
@@
Representative 12 6 6
Non-representative 1 9 6 3
Non-representative 2 12 9 3
@@
* 3 B labels duplicated
Table 2: Four training datasets for each bias value ($\beta$)
We then create two unbalanced datasets. Non-representative 1 randomly deletes three B labels for each tweet from the Representative Dataset. Non-representative 2 adds three additional A labels, drawn from $p_{i}^{A}$, to the Non-representative 1 dataset. The Non-representative 2 Dataset is more unbalanced than Non-representative 1, but contains the same number of annotations as the Representative dataset.
Finally, we use the PAIR algorithm to create the Adjusted Dataset. It is the same as the Non-representative 1 dataset, but the three B annotations are duplicated. This duplication is an easy way to adjust the unbalanced training dataset to reflect the population. Appendix §A provides a general version of the PAIR algorithm which can handle imbalances across multiple annotator characteristics. Figure 2 shows the percentage of tweets labeled OL in the four datasets for each value of $\beta$.
Figure 2: Percentage of tweets annotated as OL, by dataset and bias ($\beta$)
3.2 Model Training and Evaluation
@@
We train models on each dataset. We divide each dataset, at the tweet level, into training (2000 tweets), development (500), and test (500) sets. Each tweet appears 12 times in the Representative, Non-representative 2, and Adjusted data sets and nine times in the Non-representative 1 set.
@@
We used RoBERTa base Liu et al. (2019) as our text classifier, training for five epochs on each dataset, with development set optimization. To ensure reliable results, we trained five versions with different random seeds and averaged their performance. Appendix §B contains details on model training.
@@
We evaluate models using both calibration and accuracy metrics on the test set. While accuracy metrics directly measure classification performance, calibration metrics provide crucial insights into model reliability by assessing probability estimate quality – particularly important for high-stakes applications requiring trustworthy confidence measures.
For calibration, we report Absolute Calibration Bias (ACB), which measures how well model predicted probabilities match true frequencies in the Representative dataset (lower is better).
@@
For accuracy, we report the F1 score.
@@ 4 Results
Calibration.
Figure 3(1) compares the ACB in the test set for models trained on the simulated datasets. The dark lines show average ACB across the five training runs and the shading shows the standard deviation.
The ACB for the models trained on the Adjusted dataset closely tracks that for the Representative (gold standard) data set and does not increase with $\beta$. ACB for the models trained on the two unbalanced datasets is greater and grows with $\beta$. These results demonstrate the effectiveness of our adjustment method. Duplicating the labels from the underrepresented annotator type to match population proportions improves model calibration.
Figure 3: Metrics for OL models, by dataset and bias ($\beta$)
Accuracy.
Figure 3(2) compares the models’ F1 scores. In contrast to Figure 3(1), we do not see strong differences between the models trained on the different datasets. For all datasets, model performance declines with $\beta$: as the amount of bias in the labels increases, the models have a harder time predicting the binary OL label.
Because the F1 metric focuses on binary predictions, it is less sensitive to training biases than calibration metrics like ACB, which more explicitly capture biases through prediction scores. These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models: binary classification metrics can obscure such effects. In decision-making, miscalibrated predictions can have harmful consequences when, for example, hateful content remains undetected Van Calster et al. (2019).
Ablation Study 1: Hate Speech labels.
Appendix §C gives the ACB and F1 results for the HS labels. The PAIR approach has limited effectiveness when label distributions are highly skewed. This failure is likely due to the rarity of hate speech labels in the dataset (only 16.7% positive labels) combined with our simulation setup, which further reduces positive labels in the unbalanced datasets. PAIR may need modifications, such as stratified sampling or adaptive replication, when working with highly imbalanced datasets.
Ablation Study 2: Difficult Tweets.
A representative pool of annotators may be more important for instances that are difficult or about which people disagree. Filtering our tweets in this way also removes cases where floor and ceiling effects could mask the impact of annotator characteristics. To test PAIR performance on difficult tweets, we reran the models on tweets where $0.60.4<p_{i,HS}<$ 0.6, $0.60.4<p_{i,OL}<$ 0.6. Performance on OL is still good, and performance on HS is improved (Appendix §D).
@@ 5 Discussion & Recommendations
Our results show that (RQ1) OL prediction models perform less well when trained on data from non-representative annotator pools, and (RQ2) simple statistical adjustments can improve model calibration without collecting additional annotations. These findings establish a promising bridge between survey statistics and machine learning - offering a practical approach to make AI systems more representative of and responsive to the populations they serve, particularly for tasks involving subjective human judgments.
We recommend these four steps to reduce bias due to non-representative annotator pools:
  1. 1)
    Use social science research to identify the annotator characteristics that influence the propensity to engage in annotation and the annotations provided Eckman et al. (2024).
  2. 2)
    Collect these characteristics from annotators and gathering corresponding population-level data from national censuses or high-quality surveys. [^5] Collection and release of annotator characteristics or weights derived from them may raise confidentiality concerns. The survey literature contains relevant approaches (see Karr, 2016, for a review). Collecting annotator characteristics may also require involvement of Institutional Review Boards or other participant protection organizations Kaushik et al. (2024).
  3. 3)
    Calculate weights that match the annotators to the population on those characteristics (Bethlehem et al., 2011; Valliant et al., 2013).
  4. 4)
    Use these weights in model training. Our simple duplication approach showed promise, future work should test more sophisticated weighting approaches.
@@
Stylized Biases and Simulated Data.
@@
We have created only one version the four datasets for each label type and value of $\beta$, each of which contains random draws from the Bernoulli distribution. A more traditional statistical approach would create multiple versions of the datasets and train models on each one, to average over the sampling variability. We have not done that in this preliminary study because of the high cost and time needed to fine tune many RoBERTa models. As discussed, we used five seeds in model training.
Need for Population Benchmarks and Annotator Characteristics.
PAIR requires high quality benchmark information about the relevant population. These benchmarks might come from national statistical offices or national surveys. Annotators must provide accurate data on the same characteristics available in the benchmark data. In addition, theory from survey science demonstrates that bias will be reduced only when the characteristics used in weighting correlate with the annotations (Eckman et al., 2024). In our simulation, differences in annotations were driven solely by group membership ($A$, $B$). In the real world, it is challenging to know what characteristics impact annotation behavior and to find good benchmarks for those characteristics.
@@
While we measured calibration and accuracy, we did not examine other important metrics such as fairness across subgroups or robustness to adversarial examples. Future work on training data adjustment should assess a broader range of performance measures.
@@ Acknowledgments
Appendix A PAIR Algorithm Implementation Details
The adjustment we use to make the annotator pool more representative of the target population is a form of pseudo-population generation (Quatember, 2015). We create the pseudo-population by first constructing post-stratification weights, performing weight normalization to ensure the sum of the weights equals the size of the target population, and then duplicating each observation proportionally to its weight via deterministic replication.
Post-stratification is a method of statistical adjustment that makes a selected sample more closely resemble a target population (Bethlehem et al., 2011; Valliant et al., 2013). We form strata (groups) of annotators. For example, we might form 6 strata from three categories of region and two categories of age. Post-stratification requires population-level totals or proportions for each stratum and corresponding case-level observations. The weight for each unit $i$ in stratum $s$ is:
$w_{i}=\frac{P_{s}}{S_{s}}$ (4)
where $P_{s}$ is the true population proportion (or total) for stratum $s$ and $S_{s}$ is the sample proportion (or total) for stratum $s$. In our case, the strata of interest was a single variable (annotator Types A & B). However, post-stratification can involve multiple variables if their joint distribution is known at the population level.
where $T$ is the target total. Since we want the Adjusted dataset to match the size of our gold standard Representative dataset, the target total for the simulation was 12 labels per tweet.
Lastly, to construct a pseudo-population from our weighted data, we perform deterministic replication by replicating each unit $n_{a}$ times where:
$n_{i}=\text{round(}w_{i}^{\text{normalized}}\text{)}$ (6)
Rounding the normalized weights ensures that $n_{a}$ are integers and prevents fractional replicates.
In this initial work, we prioritized this approach for its simple interpretation and reproducibility. However, researchers may prefer other approaches, such as replication via resampling, if they are interested in how the adjustment varies across samples, or incorporating weights directly into the loss function if they want to avoid the normalization and rounding steps.
In our simulation, after post-stratification and weight normalization, each Type A label in the Non-representative 1 dataset receives a weight of 1 and each Type B label receives a weight of 2. This resulted in an Adjusted dataset where each of the Type A labels stays the same and each of the Type B labels is duplicated once.
Appendix B Model Training Details